# Neural Post-Einsteinian Test of General Relativity with the Third Gravitational-Wave Transient Catalog

### v1

@@ Abstract
Gravitational waves (GWs) from compact binaries are excellent probes of gravity in the strong- and dynamical-field regime. We report a test of general relativity (GR) with the third GW Transient Catalog (GWTC-3) using the recently developed neural post-Einsteinian framework, both on individual events and at the population level through hierarchical modeling. We find no significant violation of GR and place a constraint that, for the first time, efficiently covers non-GR theories characterized by not only post-Newtonian deviations but also those beyond under the same theory-agnostic framework.
@@
Introduction. Having passed all experimental tests in the Solar System [1] and with binary pulsars [2], Einstein’s general relativity (GR) remains our best theory for describing gravity. Despite that, Einstein’s theory is thought to be challenged by certain theoretical issues, such as the ubiquity of singularities [3, 4] and its incompatibility with quantum mechanics [5]. Moreover, GR also struggles to explain certain observed phenomena without the inclusion of additional dark fields or a cosmological constant, such as the rotation curves of the galaxies [6, 7] and the late-time acceleration of the expansion rate of the universe [8, 9]. Considerable attention has been devoted to developing modified gravity theories, and the recent observation of gravitational waves (GWs) from compact binaries [10, 11, 12, 13, 14, 15, 16, 17, 18] has opened up a new window for testing GR against these theories in the dynamical- and strong-field regime [19].
@@
Given the numerous proposals for modified gravity and the computational demand of GW model building and data analysis, GW tests of GR benefit significantly from a theory-agnostic method, since the latter can lead to robust and efficient inferences about the nature of gravity. One of the first examples of such a theory-agnostic formalism is the parametrized post-Einsteinian (ppE) framework [20, 21, 22, 23, 24, 25], which constructs a meta-model for small deviations from GR through waveform amplitude and phase corrections. In the inspiral of compact binaries, the latter is prescribed through a post-Newtonian (PN) expansion [^1] The PN formalism expands inspiral quantities in powers of $v/c$, where $v$ is the orbital velocity and $c$ is the speed of light. The expansion can be further cast into powers of GW frequency $f$ through the PN version of Kepler’s third law. [26], with the introduction of ppE theory parameters that control the type and the magnitude of the deviation. Such a formalism has been implemented successfully both by the LVK (in their “parametrized inspiral test of GR” [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]), as well as by several other groups [21, 23, 38, 39, 40, 41], in parameter estimation and model selection using both synthetic and real GW data. In all such studies, a subset of the theory parameters (i.e. those that control the type of GR deviation) is held constant, and parameter estimation is carried out only on the remaining ppE parameters (i.e. those that control the magnitude of the GR deviation).
@@
The ppE approach is theory-agnostic because a broad class of modifications to GR can be mapped to PN dephasings during the inspiral [20, 21, 22, 23, 24, 25], as long as the strength of the modification is reasonably small. For example, scalar Gauss–Bonnet (sGB) gravity [42, 43] and dynamical Chern–Simons (dCS) gravity [44, 45] contribute to a dephasing that starts at 1-1PN order and $2$ PN order, respectively. PpE tests (including the LVK implementation) examine GR against many theories within the above class of meta-models, taking into account only their leading PN-order contribution to the inspiral GW phase (and/or amplitude). So far, no deviations from GR have been reported [31, 33, 34, 35, 36, 16, 38, 39, 46, 47], and thus, constraints on certain non-GR theories can be extracted by mapping the posteriors of ppE magnitude parameters back to theory-specific parameters (e.g. coupling constants) (see [38] for a set of examples).
@@
Despite these constraints, the non-rejection of GR by such tests does not necessarily mean that current GW data are compatible with all possible modifications to GR that could affect the inspiral. The PN description adopted by the ppE formalism presumes that the underlying non-GR effect admits a legitimate expansion in powers of the orbital velocity or the GW frequency. However, several counterexamples have recently been discovered, like the binary inspirals of compact objects that source massive scalar fields [48, 49, 50, 51] or that have dark-photon interactions [52, 53]. In these theory-specific, non-GR examples, corrections to GR activate “suddenly,” and thus, they cannot be represented by a simple power law in frequency. Although a PN-based model may potentially detect deviations of this type [54], the recovery of the signal would be far from ideal, the strength of the test would be weakened, and its result would be biased to favor GR unless the signal-to-noise ratio (SNR) is unusually high or the deviation itself is unusually strong.
@@
Recently, a neural post-Einsteinian (npE) waveform model [55] was developed to mitigate the above problems for theory-agnostic inspiral tests of GR. Through deep-learning of a variational autoencoder [56], the npE model constructs a continuous latent space that maps dephasings from several discrete PN models. Crucially, the npE model maps non-PN dephasings to non-PN regions of the same continuous latent space. Additionally, this model improves the detection of PN deviations with higher PN-order corrections, and allows for a more efficient parameter estimation scheme, relative to prior implementations.
@@
In this work, we report the first data analysis application of the npE model to test GR with the third Gravitational Wave Transient Catalog (GWTC-3) [15]. We choose to focus on binary black hole (BBH) signals, which compose the majority of the events in the catalog, and which can be used to detect deviations in a large class of modified gravity theories, including sGB gravity, dCS gravity, Einstein-æther (EA) theory [57, 58], khronometric gravity [59, 60], non-commutative gravity [61], varying-$G$ theories [62, 63], and theories involving massive fields, such as massive sGB gravity [64, 51], which have been overlooked by previous tests. Hereafter, we use geometric units $G=1=c$.
@@
where $\tilde{h}_{\rm GR}$ is the GR waveform [^2] Most ppE tests, and the npE formalism, have focused on the dominant $(2,2)$ harmonic of the GW signal; other harmonics are related through a simple scaling [22, 65, 37]., which depends on source parameters $\vec{\Xi}$, such as binary masses and spins. Similar to the ppE meta-model, the npE waveform introduces a dephasing function $\delta\Psi_{\rm npE}$ that additionally depends on non-GR, phenomenological parameters $\vec{\zeta}$ to capture non-GR deviations. The later are modeled through a carefully-designed and tested, variational autoencoder, as described in [55]. The difference between the ppE and the npE model is that, in the latter, the dephasing function and the parametrization are “deeply learned” (in a physics-informed way) to unify and extend the PN representation of the dephasing that the ppE formalism is based on.
We follow the npE prescription of [55], which uses a two-dimensional npE parameter space $\vec{\zeta}=(\zeta_{1},\zeta_{2})$. The npE dephasing is designed to be proportional to the polar radius and antisymmetric under $\vec{\zeta}\rightarrow-\vec{\zeta}$. In polar coordinates $(\zeta_{b},\varphi)$, where the “radius” $\zeta_{b}$ can be negative and the angle $\varphi$ ranges within $0,\pi)$ accordingly, the npE dephasing model is constructed as
@@
To implement Eq. (2), we adopt the angular function $\psi(\varphi)$ developed in [55], which is learned by a variational autoencoder using a training set of leading PN-order dephasings, ranging from 4-4PN to 22PN order. This results in an ordered, quasi-equally spaced distribution of PN lines in a continuous angular region (and its sign-flipped image) in the $\vec{\zeta}$ space. We refer to the above region as “the PN region,” and its complement in the $\vec{\zeta}$ space as “the non-PN region.” This nomenclature is supported by the detailed parameter estimation results of [55] using simulated non-GR signals, where indeed the PN region captures dephasings that arise from a convergent PN series, and the non-PN region captures dephasings that cannot be represented as a simpler PN expansion (including non-smooth GR deviations). The latter also allows the npE test to examine theories that have been overlooked by previous tests, such as sGB theory with a massive scalar field. Following [55], we set $\varphi=0$ at a place where $\psi$ varies most rapidly with respect to $\varphi$, i.e. in the middle of the non-PN region.
With $\psi(\varphi)$ determined, $\kappa(\varphi)$ is then chosen to be a positive factor in the npE model to significantly break the GR prediction outside of the unit circle $|\vec{\zeta}|=1$, so that the latter can be conveniently taken as a prior boundary for a Bayesian test of GR. The motivation behind such a prior boundary is two-fold. From a theoretical perspective, many non-GR predictions are made based on a perturbative framework, which fails when the deviation from GR becomes too large. From an observational perspective, the current identification of a GW signal in the detector strain relies on the assumption that the signal roughly follows the predictions of GR, and a recovery model that deviates too much from GR can risk misidentifying noise artifacts as GR deviations.
In this work, $\kappa(\varphi)$ is learned by another neural network to approximate and interpolate the following npE prior boundary at fixed PN angles (defined by the already-learned $\psi(\varphi)$ function):
@@
is a measure inspired by effective cycles [66], which estimate the number of GW cycles incurred by the phasing function $\Psi$ as weighted by the noise power spectral density $S_{n}$. Here, we choose $S_{n}$ as an average estimate during LVK’s third observing run [67] [^3] https://dcc.ligo.org/LIGO-T2000012/public, from which most data for our test is collected. Note that the above prescription for $\kappa$ is similar to but not exactly the same as that in [55], and we elaborate more on the difference in the Supplemental Material.
Gravitational wave parameter estimation. We use LVK open data [68, 69] and focus on events selected for the LVK parametrized inspiral tests of GR [35, 36], each of which is (i) detected by at least two detectors, (ii) has a false-alarm rate less than $10^{-3}\,{\rm yr}^{-1}$, and (iii) accumulates an SNR greater than $6$ during the inspiral. We further filter the list by requiring that the sources have been confirmed as BBHs. This leaves us with 25 events (see Supplemental Material for a full list).
@@
For each event, we repeat the above parameter estimation for the strain data from each individual detector. We find 3 events for which the individual-detector posteriors appear to be incompatible with each other, implying that the npE tests on these events are likely impacted by detector-specific noise artifacts. One of the 3 events affected is GW200129_065458, which is known to have a glitch-removal artifact in the LVK open data that significantly impacted the inference of spin precession assuming GR [80]. The method of cross-checking posteriors obtained with different networks for the same event has been useful in identifying anomalies associated with noise features (see, e.g. [81, 82, 80].) We exclude the above 3 events from the results presented hereafter.
@@
To make full use of the remaining events in the catalog, we introduce a hierarchical model to combine the marginalized $\vec{\zeta}$ posteriors from individual-event npE tests. Similar to the hierarchical model [83, 84, 85] employed in the LVK analysis [35, 36], we assume that the bilateral deviation $\zeta_{b}$ follows a Gaussian distribution with a mean $\mu$ and a standard deviation $\sigma$. On the other hand, the polar angle $\varphi$ represents the type of theory deviation, and thus, it can be kept constant, as it must be shared across all events (i.e. we assume that any modification to GR impacts all BBH events detected and not only a subset).
@@
For the hierarchical inference, we choose a prior uniform over $\mu\in[-1,1]$, $\sigma\in[0,1]$, and $\varphi\in[0,\pi]$. We customize Bilby to sample over these parameters, where we reuse the individual-event npE posterior samples to compute the hierarchical likelihood. Once the hyperparameters $\{\mu,\sigma,\varphi\}$ are estimated, we extract the posterior quantile for GR ($\mu=0=\sigma$) and reconstruct the $\vec{\zeta}$ population for comparison with LVK test results. See the Supplemental Material for the detailed settings of our individual-event parameter estimation and hierarchical inference.
@@
Constraints on non-GR deviations. Figure 1 presents the individual-event marginalized posterior of $|\zeta_{b}|$ and $\varphi$. The $|\zeta_{b}|$ results suggest that all events are compatible with GR. Furthermore, from the $\varphi$ plot, we see no preference towards the non-PN region, which complements previous conclusions drawn by the ppE LVK test. Within the PN region, most events select $\varphi$ values towards the 22PN end. This is because the positive PN-order deviations are more correlated with existing GR parameters, such as the spins and mass ratio, whose effects also take place at positive PN orders. In addition, observe that a few events strongly prefer the 0PN angle, which is due to the correlation with the chirp mass (as first found in [21]), which accounts for the dominating effect of quadrupole radiation.
@@
Figure 1: Individual-event marginalized posteriors of the npE deviation $|\zeta_{b}|$ (upper panel) and the theory angle $\varphi$ (lower panel). Observe that all $|\zeta_{b}|$ posteriors are attached with GR at $\zeta_{b}=0$, suggesting no significant deviation from GR in general. For $\varphi$, gray dotted lines are added to label the directions of PN dephasings. Observe that the $\varphi$ posteriors mostly peak around the 0PN direction and the 22PN direction, which reflects the expected correlation between the npE parameters and the GR parameters through their mutual contribution to the GW phase at these PN orders. Apart from that, no significant preference is found towards those non-PN theory angles (to the left of the 4-4PN line or to the right of the 22PN line).
@@
Figure 2 presents the reconstructed $\vec{\zeta}$ population from the hierarchical npE test. Observe that the population distribution dies off well within the npE prior boundary $|\vec{\zeta}|=1$, which justifies our reuse of the individual-event posteriors for the hierarchical estimation, despite the fact that a few individual-event posteriors (e.g. the GW190512_180714 posterior in Fig. 1) do not die off sufficiently fast at the same boundary.
@@
Figure 2: Combined npE constraint using a hierarchical model. The blue contours enclose the 50% and 90% credible regions of a $\vec{\zeta}$ distribution reconstructed from the posterior of the hierarchical inference. The gray dotted lines mark special directions as annotated. Apart from the same PN lines in Fig. 1, we also show the angles where $\varphi$ is defined to be 0 and $\pi$. For comparison, we take the LVK posteriors published in [34, 35, 36] and overlay their 90% credible intervals in the npE parameter space as orange lines, whenever the mapping the applicable. Observe that our combined npE constraint is compatible with GR and roughly reproduces the LVK results. Moreover, the npE constraint suggests no significant deviation from GR in the non-PN region, as well as the area “between” integer and half-integer PN orders, where higher PN-order corrections to GR deviations reside [55].
@@
The combined npE constraint on non-GR deviations is more straightforwardly given by the 90% credible contour extracted from the hierarchically reconstructed $\vec{\zeta}$ population. Similar to the observation from the individual-event posteriors, we find no evidence for non-PN deviations. The constraint loosens near the 22PN line, approaching $|\vec{\zeta}|\lesssim 0.4$ at maximum. Moreover, the underlying estimation of the hyperparameters $(\mu,\sigma,\varphi)$ suggests a GR quantile of $Q_{\rm GR}=0.34$, i.e. GR ($\mu=0=\sigma$) cannot be rejected unless the measurement of the hyperparameters is restricted to a posterior region of credible level $34\%$.
@@
To compare our constraint with the LVK’s, we take the GWTC-3 combined posteriors of LVK ppE deviations along 1-1PN, 0PN, and positive PN orders [36, 86, 87, 88] and map their 90% credible intervals to the npE parameter space. Observe that our hierarchical npE 90% credible contour roughly reproduces the LVK results whenever the latter is within the npE PN range, although the npE constraint tends to be more conservative due to internal correlations [55] and the fact that the LVK analysis considered more events including a couple of neutron star-black holes. Unlike standard ppE tests, including the LVK’s, the npE constraint covers the non-PN region, as well as the area “between” integer and half-integer PN orders, where higher PN-order corrections to GR deviations reside [55]. Thus, the npE constraint extends the LVK results and leads to a more robust test of GR.
@@
Discussion and future prospects. We have conducted the first npE test of GR using inspirals of GWTC-3 BBHs, where we investigate deviations from GR under a theory-agnostic parametrization for both individual events and the combined population across the catalog using a customized hierarchical model. We find that the data does not support any significant deviation from GR, and thus, we place the first constraint on non-GR deviations covered within the npE parameter space. These deviations include PN dephasings from the GR signal with leading PN orders ranging from 4-4PN to 22PN, covering a broad class of theories that include dCS gravity, sGB gravity, EA gravity, khronometric gravity, non-commutative gravity, and varying-$G$ gravity.
@@
The current npE waveform model is built on neural networks trained with BBH signals, and in this work, we only (conservatively) apply the npE test to BBH events. This means we cannot make any inferences on non-GR theories that do not modify BBH signals, such as scalar-tensor theories, like Brans–Dicke theory [90, 91] and theories with dark-photon interactions in the hidden sector [52, 53]. On the other hand, these theories may leave imprints when the binary involves at least one NS, and there is an ongoing effort to upgrade the npE model so that it can be effectively applied to NS binaries [92]. Since BBHs compose the majority of the GWTC-3 sources, our results use the most information available, while still covering possible deviations that arise from wide set of theories.

### v2

@@ Abstract
Gravitational waves (GWs) from compact binaries are excellent probes of gravity in the strong- and dynamical-field regimes. We report a test of general relativity (GR) with the third GW Transient Catalog (GWTC-3) plus a few O4 events using the recently developed neural post-Einsteinian framework, both on individual events and at the population level through hierarchical modeling. We find no significant violation of GR and place a constraint that, for the first time, efficiently covers non-GR theories characterized by not only post-Newtonian deviations but also those beyond under the same theory-agnostic framework.
@@
Introduction. Having passed all experimental tests in the Solar System [1] and with binary pulsars [2], Einstein’s general relativity (GR) remains our best theory for describing gravity. Despite that, Einstein’s theory is thought to be challenged by certain theoretical issues such as the ubiquity of singularities [3, 4] and its incompatibility with quantum mechanics [5]. Moreover, GR also struggles to explain certain observed phenomena, such as the rotation curves of the galaxies [6, 7] and the late-time acceleration of the expansion rate of the universe [8, 9], without the inclusion of additional dark fields or a cosmological constant. Considerable attention has been devoted to developing modified gravity theories, and the recent observation of gravitational waves (GWs) from compact binaries [10, 11, 12, 13, 14, 15, 16, 17, 18] has opened up a new window for testing GR against these theories in the dynamical- and strong-field regimes [19].
@@
Given the numerous proposals for modified gravity and the computational demand of GW model building and data analysis, GW tests of GR benefit significantly from a theory-agnostic method, since the latter can lead to robust and efficient inferences about the nature of gravity. One of the first examples of such a theory-agnostic formalism is the parametrized post-Einsteinian (ppE) framework [20, 21, 22, 23, 24, 25], which constructs a metamodel for small deviations from GR through waveform amplitude and phase corrections. In the inspiral of compact binaries, the latter is prescribed through a post-Newtonian (PN) expansion [^1] The PN formalism expands inspiral quantities in powers of $v/c$, where $v$ is the orbital velocity and $c$ is the speed of light. The expansion can be further cast into powers of GW frequency $f$ through the PN version of Kepler’s third law. [27] with the introduction of ppE theory parameters that control the type and the magnitude of the deviation. Such a formalism has been implemented successfully both by the LIGO-Virgo-KAGRA (LVK) collaboration (in their “parametrized inspiral test of GR” [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]), as well as by several other groups [21, 23, 39, 40, 41, 42], in parameter estimation and model selection using both synthetic and real GW data. In all such studies, a subset of the theory parameters (i.e., those that control the type of GR deviation) is held constant, and parameter estimation is carried out only on the remaining ppE parameters (i.e., those that control the magnitude of the GR deviation).
@@
The ppE approach is theory agnostic because a broad class of modifications to GR can be mapped to PN dephasings during the inspiral [20, 21, 22, 23, 24, 25], as long as the strength of the modification is reasonably small. For example, scalar Gauss-Bonnet (sGB) gravity [43, 44] and dynamical Chern-Simons (dCS) gravity [45, 46] contribute to a dephasing that starts at 1-1PN order and $2$ PN order, respectively. PpE tests (including the LVK implementation) examine GR against many theories within the above class of metamodels, taking into account only their leading PN-order contribution to the inspiral GW phase (and/or amplitude). So far, no deviations from GR have been reported [32, 34, 35, 36, 37, 16, 39, 40, 47, 48, 49], and thus, constraints on certain non-GR theories can be extracted by mapping the posteriors of ppE magnitude parameters back to theory-specific parameters (e.g., coupling constants) (see [39] for a set of examples).
@@
Despite these constraints, the nonrejection of GR by such tests does not necessarily mean that current GW data are compatible with all possible modifications to GR that could affect the inspiral. The PN description adopted by the ppE formalism presumes that the underlying non-GR effect admits a legitimate expansion in powers of the orbital velocity or the GW frequency. However, several counterexamples have recently been discovered, like the binary inspirals of compact objects that source massive scalar fields [50, 51, 52, 53] or that have dark-photon interactions [54, 55]. In these theory-specific, non-GR examples, corrections to GR activate “suddenly,” and thus, they cannot be represented by a simple power law in frequency. Although a PN-based model may potentially detect deviations of this type [56], the recovery of the signal would be far from ideal, the strength of the test would be weakened, and its result would be biased to favor GR unless the signal-to-noise ratio (SNR) were unusually high or the deviation itself unusually strong.
@@
Recently, a neural post-Einsteinian (npE) waveform model [57] was developed to mitigate the above problems for theory-agnostic inspiral tests of GR. Through deep learning of a variational autoencoder [58], the npE model constructs a continuous latent space that maps dephasings from several discrete PN models. Crucially, the npE model maps non-PN dephasings to non-PN regions of the same continuous latent space. Additionally, this model improves the detection of PN deviations with higher PN-order corrections, and allows for a more efficient parameter estimation scheme, relative to prior implementations.
@@
In this Letter, we report the first data analysis application of the npE model to test GR with the third Gravitational Wave Transient Catalog (GWTC-3) [15] and events from the O4 discovery papers [16, 49, 59, 60, 61]. We choose to focus on binary black hole (BBH) signals, which compose the majority of the events considered and which can be used to detect deviations in a large class of modified gravity theories, including sGB gravity, dCS gravity, Einstein-æther (EA) theory [62, 63], khronometric gravity [64, 65], noncommutative gravity [66], varying-$G$ theories [67, 68], and theories involving massive fields such as massive sGB gravity [69, 53], which have been overlooked by previous tests. Hereafter, we use geometric units $G=1=c$.
@@
where $\tilde{h}_{\rm GR}$ is the GR waveform [^2] Most ppE tests, and the npE formalism, have focused on the dominant $(2,2)$ harmonic of the GW signal; other harmonics are related through a simple scaling [22, 111, 38]., which depends on source parameters $\vec{\Xi}$, such as binary masses and spins. Similar to the ppE metamodel, the npE waveform introduces a dephasing function, $\delta\Psi_{\rm npE}$, that additionally depends on non-GR, phenomenological parameters $\vec{\zeta}$ to capture non-GR deviations. The latter are modeled through a carefully designed and tested variational autoencoder, as described in [57]. The difference between the ppE and the npE models is that, in the latter, the dephasing function and the parametrization are “deeply learned” (in a physics-informed way) to unify and extend the PN representation of the dephasing that the ppE formalism is based on.
We follow the npE prescription of [57], which uses a two-dimensional npE parameter space $\vec{\zeta}=(\zeta_{1},\zeta_{2})$. The npE dephasing is designed to be proportional to the polar radius and antisymmetric under $\vec{\zeta}\rightarrow-\vec{\zeta}$. In polar coordinates $(\zeta_{b},\varphi)$, where the “radius” $\zeta_{b}$, can be negative and the angle $\varphi$ ranges within $0,\pi)$ accordingly, the npE dephasing model is constructed as
@@
To implement Eq. (2), we adopt the angular function $\psi(\varphi)$ developed in [57], which is learned by a variational autoencoder using a training set of leading PN-order dephasings, ranging from 4-4PN to 22PN order. This results in an ordered, almost equally spaced distribution of PN lines in a continuous angular region (and its sign-flipped image) in the $\vec{\zeta}$ space. We refer to the above region as “the PN region,” and its complement in the $\vec{\zeta}$ space as “the non-PN region.” This nomenclature is supported by the detailed parameter estimation results of [57] using simulated non-GR signals, where indeed the PN region captures dephasings that arise from a convergent PN series, and the non-PN region captures dephasings that cannot be represented as a simpler PN expansion (including nonsmooth GR deviations). The latter also allows the npE test to examine theories that have been overlooked by previous tests, such as sGB theory with a massive scalar field. Following [57], we set $\varphi=0$ at a place where $\psi$ varies most rapidly with respect to $\varphi$, i.e., in the middle of the non-PN region.
With $\psi(\varphi)$ determined, $\kappa(\varphi)$ is then chosen to be a positive factor in the npE model to significantly break the GR prediction outside of the unit circle $|\vec{\zeta}|=1$, so that the latter can be conveniently taken as a prior boundary for a Bayesian test of GR. The motivation behind such a prior boundary is twofold. From a theoretical perspective, many non-GR predictions are made based on a perturbative framework, which fails when the deviation from GR becomes too large. From an observational perspective, the current identification of a GW signal in the detector strain relies on the assumption that the signal roughly follows the predictions of GR, and a recovery model that deviates too much from GR can risk misidentifying noise artifacts as GR deviations.
In this work, $\kappa(\varphi)$ is learned by another neural network to approximate and interpolate the following npE prior boundary at fixed PN angles [defined by the already-learned $\psi(\varphi)$ function]:
@@
is a measure inspired by effective cycles [71], which estimate the number of GW cycles incurred by the phasing function $\Psi$ as weighted by the noise power spectral density $S_{n}$. Here, we choose $S_{n}$ as an average estimate during LVK’s third observing run [72] [^3] https://dcc.ligo.org/LIGO-T2000012/public, from which most data for our test are collected. Note that the above prescription for $\kappa$ is similar to but not exactly the same as that in [57], and we elaborate more on the difference in the Supplemental Material  [^4] See Supplemental Material, which includes Refs. [112, 113, 114, 115], for details about the npE parametrization of the waveform deviation, hierarchical modeling for the combined npE test, computational settings for the data analysis, posterior results from individual GW events with the diagnosis of noise artifacts, and posterior results from the combined npE test using the hierarchical model..
Gravitational wave parameter estimation. We use LVK open data [75, 76] and focus on events selected for the LVK parametrized inspiral tests of GR [36, 37], each of which is (i) detected by at least two detectors, (ii) has a false-alarm rate (FAR) less than $10^{-3}\,{\rm yr}^{-1}$, and (iii) accumulates an SNR greater than $6$ during the inspiral. We further filter the list by requiring that the sources have been confirmed as BBHs. We additionally consider GW241011_233834 and GW250114_082203, featured as multidetector, low-FAR, and inspiral-loud BBH events from the O4 discovery papers [16, 49, 59, 60, 61]. This leaves us with 27 events (see Supplemental Material footnotemark: for a full list).
@@
For each event, we repeat the above parameter estimation for the strain data from each individual detector. We find three events for which the individual-detector posteriors appear to be incompatible with each other, implying that the npE tests on these events are likely impacted by detector-specific noise artifacts. One of the three events affected is GW200129_065458, which is known to have a glitch-removal artifact in the LVK open data that can impact the inference of spin precession [87, 88, 89, 90], eccentricity [91, 92, 93], and GR deviations [94, 95] (see [96, 89] for more on the glitch mitigation of this event). The method of cross-checking posteriors obtained with different networks for the same event has been useful in identifying anomalies associated with noise features (see, e.g., [97, 98, 88]). We exclude the above three events from the results presented hereafter.
@@
To make full use of the remaining events, we introduce a hierarchical model to combine the marginalized $\vec{\zeta}$ posteriors from individual-event npE tests. Similar to the hierarchical model [99, 100, 101] employed in the LVK analysis [36, 37], we assume that the bilateral deviation $\zeta_{b}$ follows a Gaussian distribution with a mean $\mu$ and a standard deviation $\sigma$. On the other hand, the polar angle $\varphi$ represents the type of theory deviation, and thus, it can be kept constant, as it must be shared across all events (i.e., we assume that any modification to GR impacts all BBH events detected and not only a subset). We apply our model to GWTC-3, excluding the two O4 events, so that we can compare with LVK’s test results from the same catalog. We neglect the selection effect in our model as its impact on theory-agnostic inspiral tests of GR is small given current data [102, 103].
@@
For the hierarchical inference, we choose a prior uniform over $\mu\in[-1,1]$, $\sigma\in[0,1]$, and $\varphi\in[0,\pi]$. We customize Bilby to sample over these parameters, where we reuse the individual-event npE posterior samples to compute the hierarchical likelihood. Once the hyperparameters $\{\mu,\sigma,\varphi\}$ are estimated, we extract the posterior quantile for GR ($\mu=0=\sigma$) and reconstruct the $\vec{\zeta}$ population for comparison with LVK test results. See the Supplemental Material footnotemark: for the detailed settings of our individual-event parameter estimation and hierarchical inference.
@@
Constraints on non-GR deviations. Figure 1 presents the individual-event marginalized posterior of $|\zeta_{b}|$ and $\varphi$ estimated using uniform priors over $\zeta_{b}$ and $\varphi$. The $|\zeta_{b}|$ results suggest that all events are compatible with GR, except that the support from GW191129_134029 is marginal, which may be attributed to instrument noise (see a similar case in [32]). Furthermore, from the $\varphi$ plot, we see no preference toward the non-PN region, which complements previous conclusions drawn by the ppE LVK test. Within the PN region, most events select $\varphi$ values toward the 22PN end. This is because the positive PN-order deviations are more correlated with existing GR parameters, such as the spins and mass ratio, whose effects also take place at positive PN orders. In addition, observe that a few events strongly prefer the 0PN angle, which is due to the correlation with the chirp mass (as first found in [21]), which accounts for the dominating effect of quadrupole radiation.
@@
Figure 1: Individual-event marginalized posteriors of the npE deviation $|\zeta_{b}|$ (upper panel) and the theory angle $\varphi$ (lower panel). Observe that all $|\zeta_{b}|$ posteriors are attached with GR at $\zeta_{b}=0$, suggesting no significant deviation from GR in general. For $\varphi$, gray dotted lines are added to label the directions of PN dephasings. Observe that the $\varphi$ posteriors mostly peak around the 0PN direction and the 22PN direction, which reflects the expected correlation between the npE parameters and the GR parameters through their mutual contribution to the GW phase at these PN orders. Apart from that, no significant preference is found toward those non-PN theory angles (to the left of the 4-4PN line or to the right of the 22PN line).
@@
Figure 2 presents the reconstructed $\vec{\zeta}$ population from the hierarchical npE test. Observe that the population distribution dies off well within the npE prior boundary $|\vec{\zeta}|=1$, which justifies our reuse of the individual-event posteriors for the hierarchical estimation, despite the fact that a few individual-event posteriors (e.g., the GW190512_180714 posterior in Fig. 1) do not die off sufficiently fast at the same boundary.
@@
Figure 2: Combined GWTC-3 npE constraint using a hierarchical model. The blue contours enclose the 50% and 90% credible regions of a $\vec{\zeta}$ distribution reconstructed from the posterior of the hierarchical inference. The gray dotted lines mark special directions as annotated. Apart from the same PN lines in Fig. 1, we also show the angles where $\varphi$ is defined to be 0 and $\pi$. The black “+” at the center marks GR. For comparison, we take the LVK posteriors published in [35, 36, 37] and overlay their 90% credible intervals in the npE parameter space as orange lines, whenever the mapping is applicable. Observe that our combined npE constraint is compatible with GR and roughly reproduces the LVK results. Moreover, the npE constraint suggests no significant deviation from GR in the non-PN region, as well as the area between integer and half-integer PN orders, where higher PN-order corrections to GR deviations reside [57].
@@
The combined npE constraint on non-GR deviations is more straightforwardly given by the 90% credible contour extracted from the hierarchically reconstructed $\vec{\zeta}$ population. Similar to the observation from the individual-event posteriors, we find no evidence for non-PN deviations. The constraint loosens near the 22PN line, approaching $|\vec{\zeta}|\lesssim 0.4$ at maximum. Moreover, the underlying estimation of the hyperparameters $(\mu,\sigma,\varphi)$ suggests a GR quantile of $Q_{\rm GR}=0.34$, i.e., GR falls at the $34\%$ quantile of the hyperparameters posterior.
@@
To compare our constraint with the LVK’s, we take the GWTC-3 combined posteriors of LVK ppE deviations along 1-1PN, 0PN, and positive PN orders (estimated by freeing one PN coefficient at a time) [37, 104, 105, 106] and map their 90% credible intervals to the npE parameter space. Observe that our hierarchical npE 90% credible contour roughly reproduces the LVK results whenever the latter are within the npE PN range, although the npE constraint tends to be more conservative due to internal correlations [57] and the fact that the LVK analysis considered more events including a couple of neutron star black holes. Unlike standard ppE tests, including the LVK’s, the npE constraint covers the non-PN region as well as the area “between” integer and half-integer PN orders, where higher PN-order corrections to GR deviations reside [57]. Thus, the npE constraint extends the LVK results and leads to a more robust test of GR.
@@
Discussion and future prospects. We have conducted the first npE test of GR using BBHs’ inspirals from GWTC-3 and O4 discovery papers, where we investigate deviations from GR under a theory-agnostic parametrization for both individual events and the combined population across GWTC-3 using a customized hierarchical model. We find that the data do not support any significant deviation from GR, and thus, we place the first constraint on non-GR deviations covered within the npE parameter space. These deviations include PN dephasings from the GR signal with leading PN orders ranging from 4-4PN to 22PN, covering a broad class of theories that include dCS gravity, sGB gravity, EA gravity, khronometric gravity, noncommutative gravity, and varying-$G$ gravity.
@@
The current npE waveform model is built on neural networks trained with BBH signals, and in this work, we only (conservatively) apply the npE test to BBH events. This means we cannot make any inferences on non-GR theories that do not modify BBH signals, such as scalar-tensor theories, like Brans-Dicke theory [108, 109] and theories with dark-photon interactions in the hidden sector [54, 55]. On the other hand, these theories may leave imprints when the binary involves at least one NS, and there is an ongoing effort to upgrade the npE model so that it can be effectively applied to NS binaries [110]. Since BBHs compose the majority of the GWTC-3 sources, our results use the most information available while still covering possible deviations that arise from a wide set of theories.